OpenStudy (anonymous):

Solve 3 4x-1=9 x+1. a)x=2/3 b)x=1 c)x=3/2 d)x=2

OpenStudy (precal):

 is this your problem?

OpenStudy (anonymous):

can you make it clear ,, is there a bracket or what

OpenStudy (anonymous):

Combine like terms.

OpenStudy (anonymous):

Sorry it's 3^4x-1=9x+1

OpenStudy (anonymous):

is it 3^4x-1=9^x+1

OpenStudy (anonymous):